A proton moving with a velocity of 3×10⁸ ms‐¹ perpendicularly in a magnetic field strength X T. Given that the force of the proton received is 4.8×10‐¹³ N. Find the value of X.

1
Expert's answer
2021-12-10T11:30:07-0500

We can find the magnetic field strength as follows:


"F=qvBsin\theta,""B=\dfrac{F}{qvsin\theta}=\dfrac{4.8\times10^{-13}\ N}{1.6\times10^{-19}\ C\times3\times10^8\ \dfrac{m}{s}\times sin90^{\circ}},""B=0.01\ T."